HUBBLE SPACE TELESCOPE DAILY REPORT #3105 PERIOD COVERED: 0000Z (UTC) 04/29/02 - 0000Z (UTC) 04/30/02 Daily Status Report as of 120/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Five Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.2 Completed Two Sets of NICMOS/1/2/3 8944 (Filter Wheel/Mechanisms Mini-Functional Test) The Near Infrared Camera and Multi-Object Camera (NIC1, NIC2 and NIC3) was used to perform an early engineering test to verify the aliveness, functionality, operability, and electro-mechanical calibration of the NICMOS filter wheel motors and assembly. This was the first use of the NICMOS filter wheel mechanisms since they were disabled by ground command in January, 1999. This test was designed to obviate concerns over possible deformation or breakage of the fitter wheel "soda-straw" shafts due to excess rotational drag torque and/or bending moments which may be imparted due to changes in the dewar metrology from warm-up/cool-down. No anomalous incident occurred. 1.3 Completed Two Sets of STIS/CCD 8902 (Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. No problems were reported. 1.4 Completed Four Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.5 Completed Five Sets of WF/PC-2 8935 (Cycle 10 Standard Darks) The WF/PC-2 was used to obtain dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. Over an extended period these data will also provide a monitor of radiation damage to the CCDs. The proposal completed with no reported problems. 1.6 Completed STIS/CCD 9110 (A Search for Kuiper Belt Object Satellites) The Space Telescope Imaging Spectrograph (CCD) was used to investigate whether the large number of collisions thought to have taken place in the primordial Kuiper belt suggest that many Kuiper belt objects {KBOs} could have suffered binary-forming collisions similar to that which formed the Pluto -- Charon binary. Detection of such KBO satellites would allow measurement of KBO masses, would help to understand the past collisional environment of the Kuiper belt, and would give a context to the otherwise unique-seeming formation of the Pluto -- Charon binary. The proposal completed with no reported problems. 1.7 Completed WF/PC-2 8939 (Cycle 10 Internal Monitor) The WF/PC-2 was used to calibrate the internal monitor, to be run weekly to monitor the health of the cameras. No problems were encountered. 1.8 Completed STIS/CCD/MA1 9067 (UV Detectability of Bright Quasars in the Sloan Fields) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to take MAMA spectra of approximately 30 new, bright, high-redshift quasars in each of the next three cycles. The observations completed with no reported problems. 1.9 Completed STIS/MA1 8922 (FUV-MAMA Cycle 10 Flats) The Space Telescope Imaging Spectrograph (MA1) was used to obtain FUV-MAMA observations of the STIS internal Krypton lamp to construct an FUV flat applicable to all FUV modes. No problems were encountered. 1.10 Completed Two Sets of NICMOS/1/2/3 8945 (Cooling System Monitoring) The Near Infrared Camera and Multi-Object Camera (NIC1, NIC2 and NIC3) was used 1) to measure NICMOS detector performance during the cool-down and steady state operation of the NCS. 2}. to demonstrate stability {+/-0.1K} of the NICMOS detector temperature at the optimal science operating temperature. and 3}.to demonstrate repeatability {+/-0.1 K} of NICMOS detector temperature following changes from the optimal science operating temperature. No problems were encountered. 1.11 Completed WF/PC-2 9258 (Observatory Focus Monitor) The WF/PC-2 was used to more accurately define the HST focus. The HST focus drifts slowly and shows evidence of undergoing slips of a few microns at random times. The rate of the WF/PC-2 monitoring program is insufficient to track and/or understand OTA behavior in order to request timely and appropriate Secondary Mirror corrections. There were no reported problems. 1.12 Completed STIS/CCD 8904 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 in order to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.13 Completed STIS/CCD 8900 (Performance Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to measure the baseline performance and commandability of the CCD subsystem. All exposures are internals. The observations completed normally. 1.14 Completed STIS/CCD/WF/PC-2/ACS/WFC/HRC 9574 (Focus Monitor) The Space Telescope Imaging Spectrograph (CCD), WF/PC-2 and the Advanced Camera for Surveys (WFC and HRC) were used to monitor HST focus. In principle ACS/HRC should provide superior capability for monitoring the HST focus, but since most of the HST science will shift to ACS in Cycle 11 trending observations are initiated with this program to establish the focus of WFPC2/PC1 {and hence HST} and ACS/HRC {and WFC to a lessor degree} using exposures over full orbits that are taken in parallel. No problems were reported. 1.15 Completed Four Sets of STIS/CCD 9317 (Pure Parallel Imaging Program: Cycle 10) The Space Telescope Imaging Spectrograph (CCD) was used to perform the default archival pure parallel program for STIS during cycle 10. There were no reported problems. 1.16 Completed Two Sets of ACS/WFC 9575 (Default {Archival} Pure Parallel Program) The Advanced Camera for Surveys (WFC) was used to test ACS pure parallels in POMS. There were no reported problems. 1.17 Completed Two Sets of WF/PC-2 9180 (Gamma-ray Burst Progenitors: Probing Their Environment) The WF/PC-2 was used to perform a target of opportunity observation of gamma ray burster (GRB), GRB-011121. GRB astronomy is a field maturing at a phenomenal rate. Three important new observational and theoretical discoveries, formulated over the last twelve months, allow the proposer to address new, and in many cases, more sophisticated questions than could have been posed previously. These developments: the discovery of X-ray lines in GRB 991216; the observation that N_H as deduced from X-ray afterglow are one to two orders of magnitude larger than the dust extinction inferred from optical afterglow; and the growing realization that the afterglow emission may exhibit features of dust echoes, appear to offer unexpected and new diagnostics that will directly inform us about the progenitor, the circum-progenitor material and the immediate interstellar environs. There were no reported problems. 1.18 Completed Two Sets of STIS/CCD 9088 (Next Generation Spectral Library of Stars) The Space Telescope Imaging Spectrograph (CCD) was used to produce a "Next Generation'' Spectral Library of 600 stars for use in modeling the integrated light of galaxies and clusters by using the low dispersion UV and optical gratings of STIS. The library will be roughly equally divided among four metallicities, very low {Fe/H < -1.5}, low {-1.5 < Fe/H < -0.5}, near-solar {-0.5 < Fe/H < 0.1}, and super-solar {Fe/H > 0.1}, well-sampling the entire HR-diagram in each bin. Such a library will surpass all extant compilations and have lasting archival value, well into the Next Generation Space Telescope era. No problems occurred. 1.19 Completed ACS/WFC 9029 (Grism/Prism Performance Check) The Advance Camera for Surveys (WFC) was used to observe a Wolf-Rayet star and a White Dwarf the grism of both the Wide Field and eventually the High Resolution Channels to measure: 1} the dispersion of the grism and its field dependence; 2} the grism throughput and its field dependence; 3} the frequency of the flat field variation {L-flat} as a function of wavelength; 4} the fringe pattern at longer wavelengths as a function of position on the chip. These measurements will be carried out at several positions on the chip including the centre and the corners of the WFC and HRC chips. All observations completed without incident. 1.20 Completed WF/PC-2 9318 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform the generic target version of the WFPC2 Archival Pure Parallel program. The program was used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. There were no reported problems. 1.21 Completed STIS/MA2 8923 (FUV-MAMA Cycle 10 Flats) The Space Telescope Imaging Spectrograph (MA2) was used to obtain FUV-MAMA observations of the STIS internal Krypton lamp to construct an FUV flat applicable to all FUV modes. No problems were encountered. 1.22 Completed STIS/MA2 9573 (NUV-MAMA Daily Dark Monitor) Abstract The Space Telescope Imaging Spectrograph (CCD and MA2) was used to perform daily monitoring of the NUV MAMA detector dark noise, monitoring the effects of thermal changes on the NUV dark rate. No anomalous activity was noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 8 Successful: 8 Scheduled Re-acquisitions: 5 Successful: 5 2.2 FHST Updates: Scheduled: 22 Successful: 22 2.3 Operations Notes: ROP SR-1A was used to clear the SSR-1 EDAC error counter at 119/1422Z and to reset the TRI module at 119/1912Z. Per an operations request, the gyro reconfiguration to take Gyo #1 out of the control loop and configure to a 342 gyro configuration proceeded nominally starting at 119/1612Z. Gyro #1 was removed from the control law shortly thereafter and powered off at 119/1622Z. PSEA configuration memory and gyro safing tests were reconfigured by 119/1645Z. A bias update was performed to remove any residual drift at 119/1935Z. The first Guide Star Acquisition following the reconfiguration, at 119/2020Z, was successful. Another NICMOS filter wheel test, which began at 119/1755Z as directed by an operations request, performed as expected with results closely matching results obtained in 1997. Another such test will be performed today. The ephemeris table was uplinked at 120/0004Z per ROP DF-07A. A TTR was written for a required re-transmit at 120/0846Z during a 486 load. Using ROP DF-18A, the CCS engineering status buffer limits were adjusted at 120/0847Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: The first release of Advanced Camera for Surveys (ACS) images will be released today at 1:00 p.m. in the NASA Headquarters auditorium. Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW